The length of a rectangle is 2 more than twice the width. Find the dimensions of the rectangle if the perimeter is 76.
If x represents the rectangle's width, then which expression represents the perimeter?

Answer by mccravyedwin(421)

The length of a rectangle is 2 more than twice the width.
    L = 2W + 2

Find the dimensions of the rectangle if the perimeter is 76.
    P = 2L + 2W
   76 = 2(2W + 2) + 2W 
   76 = 4W + 4 + 2W
   76 = 6W + 4
   72 = 6W
 72/6 = W
   12 = W

    L = 2W + 2
    L = 2(12) + 2
    L = 24 + 2
    L = 26

If x represents the rectangle's width, then which expression represents the perimeter?
If you're talking about the same rectangle, then the best expression for the
perimeter is P = 76. 

If you're talking about a different rectangle, with width x and not width 12,
then the best expression for the perimeter is to replace W in the formula for
the perimeter by x. P and L will also be different.  You didn't give any facts
about how the length and width are related for this new rectangle. 

    P = 2L + 2W

    P = 2L + 2x

Answer: The expression is 2L + 2x.

Edwin